if 86.1 g of nitrogen reacts with lithium, how many grams of lithium will react


1
Expert's answer
2021-05-04T14:15:40-0400

Solution:

Balanced chemical equation:

6Li + N2 → 2Li3N

According to the equation above: n(Li)/6 = n(N2)


The molar mass of N2 is 28.013 g/mol.

Hence,

n(N2) = (86.1 g N2) × (1 mol N2 / 28.013 g N2) = 3.07357 mol N2


n(Li) = 6 × n(N2) = 6 × 3.07357 mol = 18.44 mol Li

The molar mass of Li is 6.941 g/mol.

Hence,

m(Li) = (18.44 mol Li) × (6.941 g Li / 1 mol Li) = 127.992 g Li = 128 g Li

Mass of Li = 128 g


OR (short form soluion):

(86.1 g N2) × (1 mol N2 / 28.013 g N2) × (6 mol Li / 1 mol N2) × (6.941 g Li / 1 mol Li) = 128.0 g of Li


Answer: 128.0 grams of lithium (Li) will react.